OpenStudy (wcrmelissa2001):

Help with functions please :D f(x)=ax+1 g(x)=bx-1 h(x)=c(x^2)+d where a b c d are constants and ab≠0 (for all three functions, domain of x is real number) 1. If fg=gf, show that f(x)+g(x) is independent of a and b. From here I got the relationship b=2-a (i know how to do this quesiton but i just wanna check if my relation is right) 2. From the relationship between a and b that was found in 1, find the values of a,b,c and d if hg(x)=2x^2-4x-1. Answers: a=1, b=1, c=2, d=-3. I tried comparing coefficient but got d=-2

OpenStudy (welshfella):

gf = fg b(ax + 1) - 1 = a(bx - 1) + 1 bax + b - 1 = bax - a + 1 b - 1 = 1 - a b = 2 - a you are correct

OpenStudy (welshfella):

hg(x) = c((bx - 1)^2 + d) = c(b^2x^2 - 2bx + 1 ) + d = b^2c x^2 - 2bcx + c + d compare 2x^2 - 4x - 1 b^2c = 2 -2bc = -1 and c + d = -1 solve these system of equations to find b,c and d then use b = 2 - a to find a.

OpenStudy (wcrmelissa2001):

THANKS

OpenStudy (welshfella):

yw